# Moving Out: Physically-grounded Human-AI Collaboration

### v1

@@ 1 Introduction
Humans can quickly adapt their actions to physical attributes (e.g., sizes, shapes, weights, etc.) or physical constraints (e.g., moving with stronger forces, navigating narrow paths, etc) when collaborating with other agents in the physical world. This ability is critical when embodied agents (e.g., robots) need to collaborate with humans to complete real-world tasks, such as assembly, transporting items [1, 2, 3, 4, 5], cooking [6], and cleaning [7, 8]. In these scenarios, successful interactions require understanding physical attributes and constraints while aligning with human behavior.
@@
Prior work [6, 9, 10, 8, 11, 12] has explored human-AI collaboration in discrete space or task level, which often has simplified interaction dynamics compared to the real world. As shown in Fig. 1, physically grounded tasks have increased the diversity of physical constraints, physical variations, and human behavior. While physical constraints, such as narrow passages, restrict movement and require precise coordination, there are still a large number of rotations and ways of holding objects that can lead to successful collaborations. In this paper, we propose Moving Out, a novel benchmark inspired by the Moving Out game [13], to address dynamic physical interactions and diverse collaboration scenarios in a physically grounded setting.
While AI-AI collaboration can achieve strong collaborative performance through methods like self-play [14], the resulting AI agents often struggle to adapt to human-AI collaboration, where human partners exhibit diverse and unpredictable behaviors [15]. This is particularly pronounced in physically grounded settings where minor variations in human actions, such as lifting angles or applying force, can significantly affect outcomes. An agent needs to understand the physical consequences of actions to generalize behavior across different scenarios.
@@ 2 Related Work
Environments for Human-AI Collaboration  Several multi-agent environments [16, 17] have been proposed for multi-agent reinforcement learning (MARL). Many of them are competitive rather than cooperative. For studying human-AI collaboration, existing environments have focused on different aspects. OvercookedAI [6], LBF [11], and RWARE [11] are discrete environments where agents coordinate to pick and place items, but they have limited physical attributes and constraints. The discrete state-action space also limits the diversity of agent behaviors. It Takes Two [9] involves two agents carrying a table together, but since the agents are bound to the table and cannot move freely, it does not capture scenarios requiring independent task division, collision avoidance, or awareness of when help is needed. HumanTHOR [7] and Habitat 3.0 [8] provide realistic simulation, but the collaboration in these simulators are task level or coordination in navigation. Watch and Help [18] evaluates human-AI collaboration from a social perception perspective, overlapping with our awareness collaboration mode. But their agent collaboration only considers task division and does not include physical attributes. Other studies [19, 20] also explored human-AI collaboration using the Hanabi game [21], but Hanabi is based on sequential cooperation rather than simultaneous embodied teamwork. Moving Out provides diverse collaboration modes and focuses on how AI can learn to collaborate with humans to improve task performance in continuous environments with diverse physical attributes, constraints, and human behavior. For a summarized comparison, see Appx. A.
@@
Learning Human-AI Collaboration Policy  Behavior Cloning (BC) learns human-AI collaboration policies from expert data, commonly using models like MLP [22], GRU [23], or diffusion policies [24]. Building upon BC, several studies have explored simulation and action selection methods relevant to our approach in various contexts. For instance, Wang et al. [25] utilized future state prediction for scoring actions in autonomous driving. In multi-agent scenarios,  Yuan et al. [26] proposed AgentFormer for forecasting interactive agent behaviors. Concurrently, Ding et al. [27] investigated the use of world models to predict future states based on internal knowledge. Furthermore, q-VAE [28] applies latent space transformations for next-state prediction, and Zhao et al. [29] introduced a trajectory prediction method that scores entire future trajectories to inform decision-making. Additionally, some approaches leverage RL to train collaborative agents. Building upon self-play [14] and population-based training, several works [15, 30, 31, 32, 33, 34, 35, 36] enhance agents’ ability to achieve zero-shot coordination by encouraging behavioral diversity during training and exploration. In addition, multi-agent RL algorithms, like MAPPO [37] and MADDPG [38], can also be employed to train agents for potential human-AI collaboration. However, these approaches train only through self-play and do not use human data. Some studies [39, 15] further align policies with human behavior by integrating BC models trained with human data into self-play, so that human data stays in the RL learning loop.
@@
Evaluating Human-AI Collaboration  Research on human-AI collaboration has focused on evaluating and improving AI agents across different settings. Some works redefine evaluation criteria beyond task performance, incorporating aspects like trust and perceived cooperativity [20]. In AI-assisted decision-making scenarios, prior work [40] directly computes the accuracy of AI decisions. Some works [41, 31, 36] focus on training RL agents to adapt to diverse partners and evaluate the agents by the score when playing with humans. Similarly, LLM-based agents [42] are evaluated by final score and time efficiency. Several works [36, 20, 19, 43, 44] design questionnaires to evaluate different aspects like human-like, trustworthiness, fluency, and warmth.
@@ 3 Problem Definition
We model human-AI collaboration as a decentralized Markov decision process (Dec-MDP) [45, 46], defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},\gamma,T)$, where $\mathcal{S}$ is the joint state space, and $\mathcal{A}=\mathcal{A}^{i}\times\mathcal{A}^{j}$ is the joint action space of the two agents. The transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the probability of getting the next state given a current state and a joint action. The reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ specifies the reward received for each state-joint-action pair. The observation function $\mathcal{O}:\mathcal{S}\rightarrow\mathcal{O}^{i}\times\mathcal{O}^{j}$ generates an observation for each agent for a given state. The observation of each agent makes the state jointly fully observable. The discount factor $\gamma\in[0,1]$ determines the importance of future rewards, and $T$ is the time horizon of the task.
@@
Challenges when Collaborating with Humans  When one of the agents is a human, the human agent may have diverse behaviors [6]. The AI agent must adapt its policy $\pi^{i}$ to a wide range of potential human policies $\pi^{j}$. At inference time, we assume that the real human policy $\pi^{j}$ is drawn from a universal but unknown human policy distribution $\mathcal{D}$. Thus, the AI agent’s optimal policy is:
@@
These constraints create several narrow transitions, similar to prior studies about motion planning [47, 48, 49], and can further affect the agents’ collaboration strategies. For example, in scenarios where the agents need to move a rectangular sofa through a narrow doorway, the agents need to grasp the shorter sides of the sofa to move and coordinate their moves to ensure they can fit through the entrance without collision. In this paper, we study human-AI collaboration under the challenges of continuous state space and constrained transitions introduced by physical embodiments and environments.
@@ 4.1 Environment
To test how physical environments can affect human-AI collaboration, we need an environment that follows physics. We build Moving Out on top of a single-agent environment Magical [50, 51] where agents and objects are physical bodies moving in a 2D physics simulation. As shown in Fig. 2, each agent can maneuver freely in Moving Out and can move objects with varying degrees of difficulty depending on the object size and shape. The goal is to transport all objects to the goal regions. This design emphasizes flexibility, allowing agents to act independently while also creating scenarios where collaboration is necessary for efficient task completion.
@@ 4.1.2 Layout Types
The physical variables introduce diverse collaborative behavior in Moving Out as illustrated in Fig. 3. A successful collaboration usually requires a mixture of different behaviors. Specifically, we designed 12 maps that focus on three collaboration modes. See example maps in Fig. 4 and Appx. K for the full set of maps.
@@
Coordination  The maps in this category only include small items, so each agent can complete the task independently. However, we add narrow passages in the maps. These passages often block an agent’s route, requiring the other to move aside or pass the item. For example, in Map 4 (Priority Pick), the blue agent must pick up the item and, because of the narrow passage, pass it to the pink agent. This setup enforces cooperation, as the task cannot be completed without coordination between the two agents.
@@
Action Consistency  This scenario requires agents to maintain consistent and synchronized actions over time, such as continuously aligning their efforts to move and rotate large items together. The challenge is aligning force directions and dynamically adjusting them to ensure efficient movement while navigating around tight spaces or obstacles. For instance, in Map 10 (Single Rotation), two agents must collaboratively transport a large item through a series of narrow passages. Throughout this process, the agents must continually synchronize their actions to rotate and adjust the item’s angle, allowing it to fit through the openings. Misalignment in their efforts could result in the item becoming stuck or unnecessary movements that waste time and energy.
@@ 4.3 Dataset
The data collection was approved by the Institutional Review Board (IRB). Two human players control the agents with joysticks. The game ran at 10Hz, and on average, each map took around 30 seconds (or 300 time steps) to complete the transportation of all items. See Appx. G for details.
@@ 5.1 Collaboration Behavior Augmentation
Generating New States by Perturbing the Partner’s Pose For a given trajectory, we generate new states by introducing noise to the partner’s pose while keeping all other state variables unchanged. This perturbation creates additional observation variations in training data, allowing the agent to experience a broader range of possible partner behaviors. Since human actions naturally vary, this approach helps improve the agent’s robustness to small deviations in the partner’s movements while maintaining its own task objectives. This perturbation is expressed as $\tilde{p}_{\text{partner}}=p_{\text{partner}}+\epsilon,\epsilon\sim\mathcal{N}(0,\sigma^{2})$ where $p_{\text{partner}}$ is the original partner’s pose, $\epsilon$ is Gaussian noise with mean 00 and variance $\sigma^{2}$, and $\tilde{p}_{\text{partner}}$ is the perturbed pose used to generate new state variations.
@@ 6.1 Settings
  • Diffusion Policy (DP) [24] captures multimodal distribution and has demonstrated strong performance across various tasks.
  • MAPPO [37] is a commonly used multi-agent RL algorithm. It has demonstrated strong performance in cooperative games. See Appx. C for details about training and reward settings.
Evaluation Metrics  We measure the success of collaboration using the following metrics: (1) Task Completion Rate (TCR) for size-weighted successful item delivery; (2) Normalized Final Distance (NFD) for the distances between objects and the target, measuring partial progress; (3) Waiting Time (WT) for the amount of time an agent waits for assistance with large items; and 4) Action Consistency (AC) for the degree of force alignment when moving items jointly, indicating coordination efficiency. Detailed definitions are in Appx. E.
Human Subject Study  Our study was approved by the IRB. We conducted a human subject study with 32 participants to evaluate BASS against the DP baseline in both tasks. Each participant played 32 maps in total, cooperating with each method in two rounds per map. After completing the first round, the participant and model switched to control the other agent. Upon finishing all maps, participants were given a questionnaire to capture subjective feedback. See Appx. I for details.
@@ 6.2 Results
Collaboration with AI Itself (RQ1)  Fig. 6 compares BASS with baselines in Task 1 and Task 2 by running 20 times on each map with different random seeds. BASS outperforms baselines, especially in task completion metrics TCR and NFD. For WT, DP performs better than GRU, MLP, and MAPPO, implying that it more promptly assists the other agent whenever help is needed. In terms of action consistency, there is no significant gap among methods; a potential reason is that each model is paired with itself and thus pursues the same objectives. Overall, the RL agents exhibit behaviors that follow the dense reward setting, while the BC agents mainly reflect behaviors present in the dataset. A detailed comparison of the behaviors between BC and RL agents can be found in the Appx. D.
@@ 7 Conclusion
Limitations and Future Work  While we conduct user studies using models, achieving smooth human-AI collaboration with generative models (e.g., DP, LLMs) remains challenging because of low inference speed. In our study, it requires at least 5~10Hz to have a smooth interaction (see Appx. H.2.) Moving Out aims to include diverse physical variations and collaboration modes; however, we do not cover all possible physical interactions. Future work includes improving generative models’ inference speed to achieve smooth human-AI collaboration in physical environments, leveraging LLMs’ reasoning abilities in physically grounded collaboration tasks, and extending to more complex cooperation dynamics among multiple AI agents and humans.
@@ Appendix A Comparison with Other Environments
Appendix B Compare with Oracle Simulation
@@ Appendix C MAPPO Training Setting
To train MAPPO, we integrate the Moving Out environment into the BenchMARL [52] multi-agent RL library. Our approach to MAPPO training was designed to align with the objectives of Task 1 and Task 2, which were initially conceptualized with dataset-driven methods in mind. We adapted the conditions for MAPPO as follows:

### v2

@@ 1 Introduction
Humans can quickly adapt their actions to physical attributes (e.g., sizes, shapes, weights, etc.) or physical constraints (e.g., moving with stronger forces, navigating narrow paths, etc) when collaborating with other agents in the physical world. This ability is critical when embodied agents (e.g., robots) need to collaborate with humans to complete real-world tasks, such as assembly, transporting items Semeraro et al. (2023); Munzer et al. (2018); Grigore et al. (2018); Akkaladevi et al. (2019); Ahmad et al. (2020), cooking Carroll et al. (2020), and cleaning Wang et al. (2024a); Puig et al. (2023). In these scenarios, successful interactions require understanding physical attributes and constraints while aligning with human behavior.
@@
Prior work Carroll et al. (2020); Ng et al. (2022); Papoudakis et al. (2021); Puig et al. (2023); Christianos et al. (2020); Du et al. has explored human-AI collaboration in discrete space or task level, which often has simplified interaction dynamics compared to the real world. As shown in Fig. 1, physically grounded tasks have increased the diversity of physical constraints, physical variations, and human behavior. While physical constraints, such as narrow passages, restrict movement and require precise coordination, there are still a large number of rotations and ways of holding objects that can lead to successful collaborations. In this paper, we propose Moving Out, a novel benchmark inspired by the Moving Out game SMG Studio (2020), to address dynamic physical interactions and diverse collaboration scenarios in a physically grounded setting.
While AI-AI collaboration can achieve strong collaborative performance through methods like self-play Tesauro (1994), the resulting AI agents often struggle to adapt to human-AI collaboration, where human partners exhibit diverse and unpredictable behaviors Carroll et al. (2019). This is particularly pronounced in physically grounded settings where minor variations in human actions, such as lifting angles or applying force, can significantly affect outcomes. An agent needs to understand the physical consequences of actions to generalize behavior across different scenarios.
@@ 2 Related Work
Environments for Human-AI Collaboration  Several multi-agent environments Leibo et al. (2021); Terry and Black (2020) have been proposed for multi-agent reinforcement learning (MARL). Many of them are competitive rather than cooperative. For studying human-AI collaboration, existing environments have focused on different aspects. OvercookedAI Carroll et al. (2020), LBF Christianos et al. (2020), and RWARE Christianos et al. (2020) are discrete environments where agents coordinate to pick and place items, but they have limited physical attributes and constraints. The discrete state-action space also limits the diversity of agent behaviors. It Takes Two Ng et al. (2022) involves two agents carrying a table together, but since the agents are bound to the table and cannot move freely, it does not capture scenarios requiring independent task division, collision avoidance, or awareness of when help is needed. HumanTHOR Wang et al. (2024a) and Habitat 3.0 Puig et al. (2023) provide realistic simulation, but the collaboration in these simulators are task level or coordination in navigation. Watch and Help Puig et al. (2020) evaluates human-AI collaboration from a social perception perspective, overlapping with our awareness collaboration mode. But their agent collaboration only considers task division and does not include physical attributes. Other studies Siu et al. (2021); Attig et al. (2024) also explored human-AI collaboration using the Hanabi game Bard et al. (2020), but Hanabi is based on sequential cooperation rather than simultaneous embodied teamwork. Moving Out provides diverse collaboration modes and focuses on how AI can learn to collaborate with humans to improve task performance in continuous environments with diverse physical attributes, constraints, and human behavior. For a summarized comparison, see Appx. A.
@@
Learning Human-AI Collaboration Policy  Behavior Cloning (BC) learns human-AI collaboration policies from expert data, commonly using models like MLP Rumelhart et al. (1985), GRU Cho et al. (2014), or diffusion policies Chi et al. (2023). Building upon BC, several studies have explored simulation and action selection methods relevant to our approach in various contexts. For instance, Wang et al. (2024b) utilized future state prediction for scoring actions in autonomous driving. In multi-agent scenarios,  Yuan et al. (2021) proposed AgentFormer for forecasting interactive agent behaviors. Concurrently, Ding et al. (2024) investigated the use of world models to predict future states based on internal knowledge. Furthermore, q-VAE Kobayashis (2020) applies latent space transformations for next-state prediction, and Zhao et al. (2021) introduced a trajectory prediction method that scores entire future trajectories to inform decision-making. Additionally, some approaches leverage RL to train collaborative agents. Building upon self-play Tesauro (1994) and population-based training, several works Carroll et al. (2019); Jaderberg et al. (2017); Strouse et al. (2021); Yan et al. (2023); Li et al. (2023); Yu et al. (2023); Zhao et al. (2023); Sarkar et al. (2023) enhance agents’ ability to achieve zero-shot coordination by encouraging behavioral diversity during training and exploration. In addition, multi-agent RL algorithms, like MAPPO Yu et al. (2022) and MADDPG Lowe et al. (2017), can also be employed to train agents for potential human-AI collaboration. However, these approaches train only through self-play and do not use human data. Some studies Liang et al. (2024); Carroll et al. (2019) further align policies with human behavior by integrating BC models trained with human data into self-play, so that human data stays in the RL learning loop.
@@
Evaluating Human-AI Collaboration  Research on human-AI collaboration has focused on evaluating and improving AI agents across different settings. Some works redefine evaluation criteria beyond task performance, incorporating aspects like trust and perceived cooperativity Attig et al. (2024). In AI-assisted decision-making scenarios, prior work Vollmuth et al. (2023) directly computes the accuracy of AI decisions. Some works Tylkin et al. (2021); Strouse et al. (2021); Sarkar et al. (2023) focus on training RL agents to adapt to diverse partners and evaluate the agents by the score when playing with humans. Similarly, LLM-based agents Liu et al. (2023) are evaluated by final score and time efficiency. Several works Sarkar et al. (2023); Attig et al. (2024); Siu et al. (2021); McKee et al. (2024); Hoffman (2019) design questionnaires to evaluate different aspects like human-like, trustworthiness, fluency, and warmth.
@@ 3 Problem Definition
We model human-AI collaboration as a decentralized Markov decision process (Dec-MDP) Beynier et al. (2013); Boutilier (1996), defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},\gamma,T)$, where $\mathcal{S}$ is the joint state space, and $\mathcal{A}=\mathcal{A}^{i}\times\mathcal{A}^{j}$ is the joint action space of the two agents. The transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the probability of getting the next state given a current state and a joint action. The reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ specifies the reward received for each state-joint-action pair. The observation function $\mathcal{O}:\mathcal{S}\rightarrow\mathcal{O}^{i}\times\mathcal{O}^{j}$ generates an observation for each agent for a given state. The observation of each agent makes the state jointly fully observable. The discount factor $\gamma\in[0,1]$ determines the importance of future rewards, and $T$ is the time horizon of the task.
@@
Challenges when Collaborating with Humans  When one of the agents is a human, the human agent may have diverse behaviors Carroll et al. (2020). The AI agent must adapt its policy $\pi^{i}$ to a wide range of potential human policies $\pi^{j}$. At inference time, we assume that the real human policy $\pi^{j}$ is drawn from a universal but unknown human policy distribution $\mathcal{D}$. Thus, the AI agent’s optimal policy is:
@@
These constraints create several narrow transitions, similar to prior studies about motion planning Hsu et al. (2003); Saha et al. (2005); Szkandera et al. (2020), and can further affect the agents’ collaboration strategies. For example, in scenarios where the agents need to move a rectangular sofa through a narrow doorway, the agents need to grasp the shorter sides of the sofa to move and coordinate their moves to ensure they can fit through the entrance without collision. In this paper, we study human-AI collaboration under the challenges of continuous state space and constrained transitions introduced by physical embodiments and environments.
@@ 4.1 Environment
To test how physical environments can affect human-AI collaboration, we need an environment that follows physics. We build Moving Out on top of a single-agent environment Magical Toyer et al. (2020); Zakka et al. (2021) where agents and objects are physical bodies moving in a 2D physics simulation. As shown in Fig. 2, each agent can maneuver freely in Moving Out and can move objects with varying degrees of difficulty depending on the object size and shape. The goal is to transport all objects to the goal regions. This design emphasizes flexibility, allowing agents to act independently while also creating scenarios where collaboration is necessary for efficient task completion.
@@ 4.1.2 Layout Types
The physical variables introduce diverse collaborative behavior in Moving Out as illustrated in Fig. 3. A successful collaboration usually requires a mixture of different behaviors. Specifically, we designed 12 maps that focus on three collaboration modes. See example maps in Fig. 4 and Appx. L for the full set of maps.
@@
Coordination  The maps in this category only include small items, so each agent can complete the task independently. However, we add narrow passages in the maps. These passages often block an agent’s route, requiring the other to move aside or pass the item. For example, in Map 1 (Hand Off), the blue agent must pick up the item and, because of the narrow passage, pass it to the pink agent. This setup enforces cooperation, as the task cannot be completed without coordination between the two agents.
@@
Action Consistency  This scenario requires agents to maintain consistent and synchronized actions over time, such as continuously aligning their efforts to move and rotate large items together. The challenge is aligning force directions and dynamically adjusting them to ensure efficient movement while navigating around tight spaces or obstacles. For instance, in Map 12 (Sequential Rotations), two agents must collaboratively transport a large item through a series of narrow passages. Throughout this process, the agents must continually synchronize their actions to rotate and adjust the item’s angle, allowing it to fit through the openings. Misalignment in their efforts could result in the item becoming stuck or unnecessary movements that waste time and energy.
@@ 4.3 Dataset
The data collection was approved by the Institutional Review Board (IRB). Two human players control the agents with joysticks. The game ran at 10Hz, and on average, each map took around 30 seconds (or 300 time steps) to complete the transportation of all items. See Appx. H for details.
@@ 5.1 Collaboration Behavior Augmentation
Generating New States by Perturbing the Partner’s Pose For a given trajectory, we generate new states by introducing noise to the partner’s pose while keeping all other state variables unchanged. This perturbation creates additional observation variations in training data, allowing the agent to experience a broader range of possible partner behaviors. Since human actions naturally vary, this approach helps improve the agent’s robustness to small deviations in the partner’s movements while maintaining its own task objectives. This perturbation is expressed as $\tilde{p}_{\text{partner}}=p_{\text{partner}}+\epsilon,\epsilon\sim\mathcal{N}(0,\sigma^{2})$ where $p_{\text{partner}}$ is the original partner’s pose, $\epsilon$ is Gaussian noise with mean 0 and variance $\sigma^{2}$, and $\tilde{p}_{\text{partner}}$ is the perturbed pose used to generate new state variations.
@@ 6.1 Settings
  • Diffusion Policy (DP) Chi et al. (2023) captures multimodal distribution and has demonstrated strong performance across various tasks.
  • MAPPO Yu et al. (2022) is a commonly used multi-agent RL algorithm. It has demonstrated strong performance in cooperative games. See Appx. C for details about training and reward settings.
Evaluation Metrics  We measure the success of collaboration using the following metrics: (1) Task Completion Rate (TCR) for size-weighted successful item delivery; (2) Normalized Final Distance (NFD) for the distances between objects and the target, measuring partial progress; (3) Waiting Time (WT) for the amount of time an agent waits for assistance with large items; and 4) Action Consistency (AC) for the degree of force alignment when moving items jointly, indicating coordination efficiency. Detailed definitions are in Appx. F.
Human Subject Study  Our study was approved by the IRB. We conducted a human subject study with 32 participants to evaluate BASS against the DP baseline in both tasks. Each participant played 32 maps in total, cooperating with each method in two rounds per map. After completing the first round, the participant and model switched to control the other agent. Upon finishing all maps, participants were given a questionnaire to capture subjective feedback. See Appx. J for details.
@@ 6.2 Results
Collaboration with AI Itself (RQ1)  Fig. 6 compares BASS with baselines in Task 1 and Task 2 by running 20 times on each map with different random seeds. BASS outperforms baselines, especially in task completion metrics TCR and NFD. For WT, DP performs better than GRU, MLP, and MAPPO, implying that it more promptly assists the other agent whenever help is needed. In terms of action consistency, there is no significant gap among methods; a potential reason is that each model is paired with itself and thus pursues the same objectives. Overall, the RL agents exhibit behaviors that follow the dense reward setting, while the BC agents mainly reflect behaviors present in the dataset. A detailed comparison of the behaviors between BC and RL agents can be found in the Appx. E.
@@ 7 Conclusion
Limitations and Future Work  While we conduct user studies using models, achieving smooth human-AI collaboration with generative models (e.g., DP, LLMs) remains challenging because of low inference speed. In our study, it requires at least 5~10Hz to have a smooth interaction (see Appx. I.2.) Moving Out aims to include diverse physical variations and collaboration modes; however, we do not cover all possible physical interactions. Future work includes improving generative models’ inference speed to achieve smooth human-AI collaboration in physical environments, leveraging LLMs’ reasoning abilities in physically grounded collaboration tasks, and extending to more complex cooperation dynamics among multiple AI agents and humans.
@@ Appendix A Comparison with Other Environments
Appendix B Comparison with Oracle Simulation
@@ Appendix C MAPPO Training Setting
To train MAPPO, we integrate the Moving Out environment into the BenchMARL Bettini et al. [2024] multi-agent RL library. Our approach to MAPPO training was designed to align with the objectives of Task 1 and Task 2, which were initially conceptualized with dataset-driven methods in mind. We adapted the conditions for MAPPO as follows: